OpenStudy (anonymous):

can someone help me solve this problem ? let f(x)= (x^3+3x+8)^2 F'(x)= F'(2)=

hartnn (hartnn):

do you know the chain rule ?

OpenStudy (anonymous):

not really sure with the chain rule..

OpenStudy (anonymous):

need to learn that then to do a question of this type

OpenStudy (anonymous):

yes, that's why I'm trying to learn it here.. if that helps.

hartnn (hartnn):

let me give you a similar example, if \(f(x) = (ax^2+bx+c)^3 \) then by chain rule, \(f'(x)=3(ax^2+bx+c)^2 (ax^2+bx+c)' \\ =3(ax^2+bx+c)^2(2ax+b) \) got this ?

hartnn (hartnn):

by definition, \(f(g(x))' = f'(g(x)) \times g'(x)\)

OpenStudy (anonymous):

yeah

hartnn (hartnn):

so try once for your Q, maybe.. ?

OpenStudy (anonymous):

you dont subtract the exponent n-1?

hartnn (hartnn):

i made the exponent n-1 in the Q, ax^2+bx+c had exponent 3, in the answer i became 2, right ?

hartnn (hartnn):

for x^n, the derivative is \((x^n)'=n x^{n-1}\)

OpenStudy (anonymous):

okay... sort of there. but then the second equation next to it confused me a little.

hartnn (hartnn):

yeah, from the definition, i needed to multiply by the derivative of inner function also (g'(x)) so ,i multiplied by derivative of ax^2+bx+c which is 2ax+b

hartnn (hartnn):

in your Q, you'll multiply by (x^3+3x+8)

hartnn (hartnn):

**by the derivative of (x^3+3x+8)

OpenStudy (anonymous):

so the derivative for F'(x) would be.. (3x^2+3x+8)?

hartnn (hartnn):

no... you have to first derivate the out function, (x^2), then multiply the derivative of each term of x^3+3x+8 first tell whats the derivative of x^2

OpenStudy (anonymous):

2x

hartnn (hartnn):

so, derivative of (x^3+3x+8)^2 will be 2(x^3+3x+8) *(derivative of x^3+3x+8 ) , right ? can you differentiate, x^3+3x+8 ?

OpenStudy (anonymous):

2x^2+6x+16?

hartnn (hartnn):

no ? how u got that ?

OpenStudy (anonymous):

multiplied...

hartnn (hartnn):

derivative of x^3+3x+8 lets take it term by term derivative of x^3 =... ?

OpenStudy (anonymous):

3x

hartnn (hartnn):

no... derivative of x^n = nx^{n-1}

OpenStudy (anonymous):

3x^2

hartnn (hartnn):

derivative of 3x ?

OpenStudy (anonymous):

3

hartnn (hartnn):

derivative of 8 ?

OpenStudy (anonymous):

2? i really am so confused

hartnn (hartnn):

derivative of a constant = 0 so derivative of 8 is 0 so, derivative of x^3+3x+8 = 3x^2+3 right ?

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

i understand that better.

OpenStudy (anonymous):

now what if 2 was plugged in

hartnn (hartnn):

so, we have : derivative of (x^3+3x+8)^2 will be 2(x^3+3x+8) *(derivative of x^3+3x+8 ) , = 2(x^3+3x+8) * (3x^2+3) got this ?

OpenStudy (anonymous):

i thought the derivative of f '(x) is 3x^2+3

hartnn (hartnn):

didn't i mention to u the chain rule ?

OpenStudy (anonymous):

oh yeah, multiple use the one before and after

hartnn (hartnn):

so any doubts now ? that would be your final answer...

OpenStudy (anonymous):

nope no doubts. now what if the problem was f ' (2)= ?

hartnn (hartnn):

just plug in x=2

OpenStudy (anonymous):

i did but im still not sure what am i suppose to do.

OpenStudy (anonymous):

i plug them in but i get confused with the derivative and actually multiplying it out.

hartnn (hartnn):

2(x^3+3x+8) * (3x^2+3) put x=2 2(2^3+3*2+8) * (3*2^2+3) =... ?

OpenStudy (anonymous):

44

OpenStudy (anonymous):

thank you. thats correct

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

would it be possible to know the answer of a few problems because I would need to sign off and practice by hand but I would like to check my work and see if they're correct.